### v5

@@ 4 Structural Approach
This section develops a structural framework to evaluate mergers that captures key economic mechanisms affecting merger outcomes. With this approach, we can determine not just whether prices change, but also why. In particular, it allows us to decompose the price effect of a merger into efficiency, market power, and coordination effects.
Our first bridge demonstrates that the canonical DiD specification in Equation (1) emerges as a reduced-form of the equilibrium demand and supply conditions. This connection also clarifies why DiD identification may fail in merger settings and enables us to construct an aggregate measure of price effects analogous to the DiD coefficient, using demand and cost parameters, that is more robust than the DiD estimate. Our second bridge employs recent advances in synthetic control and synthetic DiD methods to improve control market selection, improving the precision of structural estimates.
@@ 4.1 Demand and Supply
Mergers can also affect competitive conduct. For instance, sustaining supracompetitive pricing may become easier when fewer firms remain post-merger, as monitoring becomes easier, which in turn facilitates tacit coordination. This aspect of the merger suggests conduct may shift toward monopolistic behavior following a merger. To capture this effect, we allow the conduct to change with the merger, such that $\lambda(I=0)=\lambda_{\text{pre}}\leq\lambda(I=1)=\lambda_{\text{post}}$.
@@
$P=c(Q,W,I;\boldsymbol{\theta})-\lambda(I)\times h(Q,Y,\boldsymbol{\alpha})+\vartheta.$ (6)
@@
We assume that mergers affect only the supply-side relationships but not the demand. This assumption reflects the principle that preferences remain constant following mergers, even as consumers adjust to the resulting price changes. Although firms may strategically modify their product attributes after a merger, such as route networks, service quality, or operational frequencies, we assume that the unobservable determinants of these adjustments are statistically independent of the demand and cost shocks in pricing equations. [^7] This orthogonality condition is implicit in the DiD framework. When estimating demand and cost parameters, we use only the characteristics of competing non-merging firms, excluding those of merging firms, to minimize the effects of product repositioning that may occur after the merger.
@@ 4.2 Identification
These linear specifications, combined with the first-order optimality condition from Equation (6) and the relationship $h(Q)=Q_{mt}/\alpha_{1}$, yield a system of simultaneous equations governing equilibrium outcomes in each market $m$ during period $t$:
@@
where the (nuisance) parameter $\tilde{\gamma}_{mt}:=(\theta_{1}-\lambda(I_{mt})/\alpha_{1})$ captures the economies of scale $\theta_{1}$, strategic conduct $\lambda(I_{mt})$ and the demand slope $\alpha_{1}$. [^8] As Abbring et al. (2025) point out, our structural approach traces back to Wright (1928), who introduced structural models of demand and supply, established identification using instrumental variables, and developed methods for counterfactual analysis nearly a century ago. For the ease of presentation, using the definition of $I_{mt}$ from (1), let $\tilde{\gamma}_{mt}=\tilde{\gamma}_{1}$ if $I_{mt}=1$ and $\tilde{\gamma}_{mt}=\tilde{\gamma}_{0}$ otherwise, and with a slight abuse of notation, we continue to use $\tilde{\boldsymbol{\theta}}$ to represent the cost parameters $(\theta_{0},\tilde{\gamma}_{1},\tilde{\gamma}_{0},\theta_{2},\theta_{3})^{\top}$.
@@ 4.2.1 Demand and Supply Parameters
The proof is in the Appendix B. While the demand and supply relations in (7) are written for one-dimensional instruments for expositional clarity, in our proof and when estimating, we use multiple IVs for each Equation. For example, for the airline industry, cost-shifters include potential competitor presence at origin and destination airports, legacy carrier presence at both endpoints, and market-specific operational characteristics (such as total distance, metal carriers, and connection requirements) that maintain shift supply conditions while remaining plausibly independent of route-specific demand fluctuations. Similarly, demand-shifters include the total number of destinations accessible from the origin airport (excluding merging parties’ routes) and demographic variables (population-based market size, births, deaths, net migration, total income, and per capita income) that influence passenger demand without directly affecting operational costs or strategic pricing considerations.
@@ 4.2.2 Strategic Conduct Parameters
The proof is in Appendix B, but here we provide an intuitive explanation of the results. The non-identification in Lemma 2-(1) arises from the substitutability between marginal cost and conduct parameters in determining $\tilde{\gamma}_{mt}$. For illustration, suppose we increase the strategic conduct, i.e., make firms less competitive, by $\kappa>0$ and the scale parameter $\theta_{1}$ by $\kappa/\alpha_{1}$. On the one hand, the economies of scale put downward pressure on prices. On the other hand, because firms are now less competitive, there is an upward pressure on prices. In the proof, we show that these opposing pressures on prices cancel out, i.e., the two sets of parameters $(\lambda_{\text{pre}},\lambda_{\text{post}},\theta_{1})$ and $(\kappa\lambda_{\text{pre}},\kappa\lambda_{\text{post}},\theta_{1}\kappa/\alpha_{1})$ are observationally equivalent. However, the change in the strategic conduct ($\Delta\lambda$) is identified as $\Delta\lambda=-\alpha_{1}\times(\tilde{\gamma}_{1}-\tilde{\gamma}_{0})$.
@@ 4.3 First Bridge: Structural to Quasi-Experimental
4.3.1 Aggregate Price Effect of Merger
An application of this equivalence is that researchers seeking an aggregate measure of merger effects on prices can still use our structural framework. In particular, the structural approach recovers the same interpretable coefficient while being robust to the fundamental identification challenges that may render the standard DiD estimate unreliable.
In particular, comparing Equations (1) and (9), the coefficient on the merger indicator $I_{mt}$ can be thought of as the aggregate price effect of the merger, which is:
@@ 4.3.1 Aggregate Price Effect of Merger
where the second last equality uses the fact that the demand elasticity is ${\mathcal{E}}_{d,mt}=\alpha_{1}\frac{P_{mt}}{Q_{mt}}$, the supply elasticity is ${\mathcal{E}}_{s,mt}=\frac{1}{\tilde{\gamma}_{mt}}\frac{P_{mt}}{Q_{mt}}$, and their ratio is $\tilde{\gamma}_{mt}\alpha_{1}$, and the last equality uses the definition of cost pass-through rate (Weyl and Fabinger, 2013). [^10] For more on merger effects and pass-through rates, see Jeffe and Weyl (2013) and Miller et al. (2016). Equation (10) clarifies that the aggregate effect of the merger equals the cost efficiency adjusted for the cost pass-through rate. It captures if and how merger-induced cost changes are passed to prices, given the market’s demand and supply elasticities.
Any such aggregate measure conflates the direct cost effect ($\theta_{3}$) with conduct changes embedded in $\tilde{\gamma}_{mt}$. A small coefficient could indicate either modest efficiency gains or low pass-through due to increased market power. Our structural approach resolves this ambiguity by identifying $\theta_{3}$, the actual cost efficiency, separate from the pass-through rate determined by demand and supply estimates, which includes the firms’ strategic conduct. As we demonstrate next, when mergers affect strategic conduct and alter supply elasticity, our structural framework can detect and correct the resulting biases that affect DiD design.
@@ 4.3.2 Robustness
of demand shocks $\varphi_{mt}$ and supply shocks $\vartheta_{mt}$, where the weights $A_{mt}$ and $B_{mt}$ depend on model primitives including the strategic conduct parameter embedded in $\tilde{\gamma}_{mt}$; see Equation (10). Because mergers can alter strategic conduct, they potentially change these very weights that determine how structural shocks manifest in observed prices, generating endogeneity that the DiD design cannot remedy.
@@
Airlines systematically select markets based on both demand prospects and cost considerations, violating at least one of these conditions. The DiD approach requires both these conditions to be satisfied, whereas the structural approach requires only the supply-side parallel trends condition (13) to be satisfied. Selection on demand unobservables does not threaten identification because mergers affect costs directly, not demand, and we instrument for endogenous quantities using our exclusion restrictions.
@@
When mergers increase market power, the weight $A_{mt}$ on demand shocks increases. Any positive demand shock in the post-merger period generates an amplified price response through enhanced market power. The DiD specification attributes this entire price increase to the merger dummy, confounding the direct merger effect with the interaction between demand shocks and changed conduct. This creates systematic bias that increases with both demand volatility and conduct changes.
@@ 5.1.1 Demand Shifters
Our primary demand instrument exploits variation in the total number of markets served from origin airports, excluding routes operated by merging airlines. This measure captures the option value of network access: broader destination availability increases travel flexibility, enables multi-city trips, and raises the probability that air travel meets specific transportation needs. Crucially, this market thickness effect makes air travel more attractive relative to driving, without affecting any individual airline’s operational costs. While carrier-specific networks affect marginal costs (Ciliberto et al., 2021), the aggregate count of destinations served by all carriers affects only demand through increased consumer awareness and consideration of air travel. We also include population-based market size (geometric mean).
@@ 5.1.2 Cost Shifters
We also count the number of distinct regional carriers used by a legacy carrier to serve a specific market. The use of regional airlines should be correlated with costs (through fuel consumption, crew requirements, aircraft utilization, and network management). However, it should not directly impact consumer utility, given the observed product characteristics.
@@ 5.2 Estimated Merger Effects
where $I_{m,t,k}=1$ if market $m$ is in quarter $k$ post-merger at time $t$, and 0 otherwise. This estimation recognizes that mergers may facilitate coordination among remaining competitors. As before, we continue to assume that the efficiency may take time to realize, but we assume that the firms’ strategic conduct may change immediately after the merger. We allow the conduct parameter to vary across periods as:
@@
Correspondingly, let $D^{\text{ctrl}}_{mt}$ be a binary variable equal to one if $m$ is a control market at time $t$, and zero otherwise. Similarly, let $D^{\text{treat-pre}}_{mt}$ and $D^{\text{treat-post}}_{mt}$ be equal to one if $m$ is treated and pre-merger (respectively, post-merger) at time $t$, and zero otherwise. Finally, let $\text{Transition}_{mt}$ be equal to one only when a market $m$ transitions from pre- to post-merger, and $\text{Transition}_{m,t,k}$ be equal to one when transitioning specifically into quarter $k$ post-merger. So, we keep track of three cases: $\text{Cases}=\{\text{ctrl, treat-pre, treat-post}\}$. Then, we can write the estimating price equation in first difference as
@@
The key insight is that the slope coefficient on $\Delta Q_{mt}$ differs across market types and time periods, reflecting different conduct parameters. Additionally, there is a level effect involving $Q_{m,t-1}$ when markets transition from a pre-merger to a post-merger state due to changes in conduct parameters. Each coefficient $\theta_{3,k}$ represents the price change when transitioning into quarter $k$ after the merger. The average efficiency effect weights each quarter’s contribution by the fraction of post-merger observations that experience it: $\sum_{k=1}^{8}\theta_{3,k}\times\frac{9-k}{8}$. Table 5 presents the estimation results.
@@
Change in Strategic Conduct. Panel B reveals evidence of post-merger coordination effects that vary substantially across mergers. For DL-NW, we observe a change in conduct parameter of $\Delta\lambda=-0.141$, indicating a decrease in coordination following the merger, although this estimate is statistically insignificant. This result likely reflects the limited number of control markets available for this merger, as the DL-NW transaction involved two large network carriers with extensive route overlap, leaving fewer comparable control markets.
UA-CO shows a modest increase in the conduct parameter ($\Delta\lambda=0.031$), though this estimate is also imprecise. In contrast, AA-US demonstrates a more substantial increase ($\Delta\lambda=0.107$, p-value 0.05)—a meaningful shift toward less competitive behavior among the remaining airlines, providing evidence that mergers can have coordinated effects.
@@
Aggregate Price Effects. Panel D presents the aggregate merger effects on prices, incorporating both efficiency and conduct channels. As with the previous estimates, none of the aggregate price effects is estimated precisely. DL-NW shows a price decrease of $\mathdollar 8.6$ per hundred dollars of fare, reflecting the combination of efficiency gain and reduced conduct. UA-CO exhibits a modest price increase of $\mathdollar 2.6$ per hundred dollars, while AA-US shows a larger price increase of $\mathdollar 10.6$ per hundred dollars. These effects reflect the combined impact of efficiency changes and altered competitive conduct following the merger.
The progressive deterioration in competitive conduct – from negative to positive to statistically significant positive – suggests that successive mergers may have facilitated increasingly coordinated behavior among the remaining carriers. Indeed, this escalating pattern of coordination effects was precisely the concern that led the US DOJ to initially block the AA-US merger in 2013, fearing that the merger would generate coordinated effects. While we cannot make definitive causal claims due to statistical limitations, the systematic temporal progression provides suggestive evidence that merger waves in concentrated industries may have cumulative effects on market competitiveness.

### v6

@@ 4 Structural Approach
This section develops a structural framework for evaluating mergers that captures key economic mechanisms affecting merger outcomes. With this approach, we can determine not just whether prices change, but also why. In particular, it allows us to decompose the price effect of a merger into efficiency, market power, and coordination effects.
Our first bridge demonstrates that the canonical DiD specification in Equation (1) emerges as a reduced-form of the equilibrium demand and supply conditions. This connection also clarifies why DiD identification may fail in merger settings and enables us to construct a measure of price effects analogous to the DiD coefficient, using demand and cost parameters, that is more robust than the DiD estimate. Our second bridge employs recent advances in synthetic control and synthetic DiD methods to enhance control-market selection, thereby improving the precision of structural estimates.
@@ 4.1 Demand and Supply
Mergers can also affect competitive conduct. For instance, sustaining supracompetitive pricing may become easier when fewer firms remain post-merger, as monitoring becomes easier, thereby facilitating tacit coordination. This aspect of the merger suggests conduct may shift toward monopolistic behavior. To capture this effect, we allow the conduct to change with the merger, such that $\lambda(I=0)=\lambda_{\text{pre}}\leq\lambda(I=1)=\lambda_{\text{post}}$.
@@
$P=c(Q,W,I;\boldsymbol{\theta})-\lambda(I)\times h(Q,Z;\boldsymbol{\alpha})+\vartheta.$ (6)
@@
We assume that mergers affect only supply-side relationships, not demand. This assumption reflects the principle that preferences remain constant following mergers, even as consumers adjust to the resulting price changes. Although firms may strategically modify their product attributes after a merger, such as route networks, service quality, or operational frequencies, we assume that the unobservable determinants of these adjustments are statistically independent of the demand and cost shocks in pricing equations. [^8] This orthogonality condition is implicit in the DiD framework. When estimating demand and cost parameters, we use only the characteristics of competing non-merging firms, excluding those of merging firms, to minimize the effects of product repositioning that may occur after the merger.
@@ 4.2 Identification
These linear specifications, combined with the first-order optimality condition from Equation (6) and the relationship $h(Q,Z;\boldsymbol{\alpha})=Q_{mt}/\alpha_{1}$, yield a system of simultaneous equations governing equilibrium outcomes in each market $m$ during period $t$:
@@
where the (nuisance) parameter $\tilde{\gamma}_{mt}:=(\theta_{1}-\lambda(I_{mt})/\alpha_{1})$ captures the economies of scale $\theta_{1}$, strategic conduct $\lambda(I_{mt})$ and the demand slope $\alpha_{1}$. [^9] As Abbring et al. (2025) point out, this structural approach traces back to Wright (1928), who introduced a model of demand and supply and the identification and estimation strategies nearly a century ago. For the ease of presentation, using the definition of $I_{mt}$ from (1), let $\tilde{\gamma}_{mt}=\tilde{\gamma}_{1}$ if $I_{mt}=1$ and $\tilde{\gamma}_{mt}=\tilde{\gamma}_{0}$ otherwise, and with a slight abuse of notation, we continue to use $\tilde{\boldsymbol{\theta}}$ to represent the cost parameters $(\theta_{0},\tilde{\gamma}_{1},\tilde{\gamma}_{0},\theta_{2},\theta_{3})^{\top}$.
@@ 4.2.1 Demand and Supply Parameters
The proof is in the Appendix B. While the demand and supply relations in (7) are written for one-dimensional instruments for expositional clarity, in our proof and when estimating, we use multiple IVs for each Equation. For example, in the airline industry, cost-shifters include the presence of potential competitors at origin and destination airports, the presence of legacy carriers at both endpoints, and market-specific operational characteristics (such as total distance, metal carriers, and connection requirements) that maintain shift supply conditions while remaining plausibly independent of route-specific demand fluctuations. Similarly, demand-shifters include the total number of destinations accessible from the origin airport (excluding merging parties’ routes) and demographic variables (population-based market size, births, deaths, net migration, total income, and per capita income) that influence passenger demand without directly affecting operational costs or pricing considerations.
@@ 4.2.2 Strategic Conduct Parameters
The proof is in Appendix B, but here we provide an intuitive explanation of the results. The non-identification in Lemma 2-(1) arises from the substitutability between marginal cost and conduct parameters in determining $\tilde{\gamma}_{mt}$. For illustration, suppose we increase the strategic conduct, i.e., make firms less competitive, by $\kappa>0$ and the scale parameter $\theta_{1}$ by $\kappa/\alpha_{1}$. On the one hand, the economies of scale put downward pressure on prices. On the other hand, because firms are now less competitive, there is upward pressure on prices. In the proof, we show that these opposing pressures on prices cancel out, i.e., the two sets of parameters $(\lambda_{\text{pre}},\lambda_{\text{post}},\theta_{1})$ and $(\kappa\lambda_{\text{pre}},\kappa\lambda_{\text{post}},\theta_{1}\kappa/\alpha_{1})$ are observationally equivalent. However, the change in the strategic conduct ($\Delta\lambda$) is identified as $\Delta\lambda=-\alpha_{1}\times(\tilde{\gamma}_{1}-\tilde{\gamma}_{0})$.
@@ 4.3 First Bridge: Structural to Quasi-Experimental
4.3.1 Price Effect of Merger
An application of this equivalence is that researchers seeking a measure of the (partial) effect of merger on prices can still use our structural framework. In particular, the structural approach recovers the same interpretable coefficient while being robust to the fundamental identification challenges that may render the standard DiD estimate unreliable.
In particular, comparing Equations (1) and (9), the coefficient on the merger indicator $I_{mt}$ can be thought of as the price effect of the merger, which is:
@@ 4.3.1 Price Effect of Merger
where the second last equality uses the fact that the demand elasticity is ${\mathcal{E}}_{d,mt}=\alpha_{1}\frac{P_{mt}}{Q_{mt}}$, the supply elasticity is ${\mathcal{E}}_{s,mt}=\frac{1}{\tilde{\gamma}_{mt}}\frac{P_{mt}}{Q_{mt}}$, and their ratio is $\tilde{\gamma}_{mt}\alpha_{1}$, and the last equality uses the definition of cost pass-through rate (Weyl and Fabinger, 2013). [^11] For more on merger effects and pass-through rates, see Jeffe and Weyl (2013) and Miller et al. (2016). Equation (10) clarifies that the effect of the merger that the DiD estimate is meant to capture equals the cost efficiency adjusted for the cost pass-through rate. It captures if and how merger-induced cost changes are passed through to prices, given the market’s demand and supply elasticities.
Any such measure conflates the direct cost effect ($\theta_{3}$) with conduct changes embedded in $\tilde{\gamma}_{mt}$. A small coefficient could indicate either modest efficiency gains or low pass-through due to increased market power. Our structural approach resolves this ambiguity by identifying $\theta_{3}$, the actual cost efficiency, separate from the pass-through rate, which is determined by demand and supply estimates and includes the firms’ strategic conduct. As we demonstrate next, when mergers affect strategic conduct and alter supply elasticity, our structural framework can detect and correct the resulting biases that affect DiD estimates.
@@ 4.3.2 Robustness
of demand shocks $\varphi_{mt}$ and supply shocks $\vartheta_{mt}$, where the weights $A_{mt}$ and $B_{mt}$ depend on model primitives including the strategic conduct parameter embedded in $\tilde{\gamma}_{mt}$; see Equation (10). Because mergers can alter strategic conduct, they can alter the very weights that determine how structural shocks manifest in observed prices, generating endogeneity that the DiD design cannot remedy.
@@
Airlines systematically select markets based on both demand prospects and cost considerations, violating at least one of these conditions. The DiD approach requires both these conditions to be satisfied, whereas the structural approach requires only the supply-side parallel trends condition (13) to be satisfied. Selection on the demand unobservables does not threaten identification because mergers affect costs directly, not demand, and we instrument for endogenous quantities using our exclusion restrictions.
@@
When mergers increase market power, the weight $A_{mt}$ on demand shocks increases. Any positive demand shock in the post-merger period generates an amplified price response through enhanced market power. The DiD specification attributes this entire price increase to the merger dummy, confounding the direct merger effect with the interaction between demand shocks and changed conduct. This creates systematic bias that increases with both demand volatility and changes in conduct.
@@ 5.1.1 Demand Shifters
Our primary demand instrument exploits variation in the total number of markets served from origin airports, excluding routes operated by merging airlines. This measure captures the option value of network access: broader destination availability increases travel flexibility, enables multi-city trips, and raises the probability that air travel meets specific transportation needs. Crucially, this market thickness effect makes air travel more attractive relative to driving, without affecting any individual airline’s operational costs. While carrier-specific networks affect marginal costs (Ciliberto et al., 2021), the aggregate number of destinations served by all carriers affects demand only by increasing consumer awareness and consideration of air travel. We also include population-based market size (geometric mean).
@@ 5.1.2 Cost Shifters
We also count the number of distinct regional carriers a legacy carrier uses to serve a specific market. The use of regional airlines should be correlated with costs (through fuel consumption, crew requirements, aircraft utilization, and network management). However, it should not directly impact consumer utility, given the observed product characteristics.
@@ 5.2 Estimated Merger Effects
where $I_{m,t,k}=1$ if market $m$ is in quarter $k$ post-merger at time $t$, and 0 otherwise. This estimation recognizes that mergers may facilitate coordination among remaining competitors. As before, we continue to assume that the efficiency may take time to realize, but the firms’ strategic conduct may change immediately after the merger. We allow the conduct parameter to vary across periods as:
@@
Correspondingly, let $D^{\text{ctrl}}_{mt}$ be a binary variable equal to one if $m$ is a control market at time $t$, and zero otherwise. Similarly, let $D^{\text{treat-pre}}_{mt}$ and $D^{\text{treat-post}}_{mt}$ be equal to one if $m$ is treated and pre-merger (respectively, post-merger) at time $t$, and zero otherwise. Finally, let $\text{Transition}_{mt}\in\{0,1\}$ be a binary variable equal to one only when a market $m$ transitions from pre- to post-merger, and $\text{Transition}_{m,t,k}\in\{0,1\}$ be equal to one when transitioning specifically into quarter $k$ post-merger. So, we keep track of three cases: $\text{Cases}=\{\text{ctrl, treat-pre, treat-post}\}$. Then, we can write the estimating price equation in first difference as
@@
The key insight is that the slope coefficient on $\Delta Q_{mt}$ differs across market types and time periods, reflecting different conduct parameters. Additionally, there is a level effect involving $Q_{m,t-1}$ when markets transition from a pre-merger to a post-merger state, driven by changes in conduct parameters. Each coefficient $\theta_{3,k}$ represents the price change when transitioning into quarter $k$ after the merger. The average efficiency effect weights each quarter’s contribution by the fraction of post-merger observations that experience it: $\sum_{k=1}^{8}\theta_{3,k}\times\frac{9-k}{8}$. Table 5 presents the estimation results.
@@
Change in Strategic Conduct. Panel B reveals evidence of post-merger coordination effects that vary substantially across mergers. For DL-NW, we observe a change in conduct parameter of $\Delta\lambda=-0.141$, indicating a decrease in coordination following the merger, although this estimate is statistically insignificant. This result likely reflects the limited number of control markets available for this merger, as the DL-NW transaction involved two large network carriers with extensive route overlap, leaving fewer control markets.
UA-CO shows a modest increase in the conduct parameter ($\Delta\lambda=0.031$), though this estimate is also imprecise. In contrast, AA-US shows a more substantial increase ($\Delta\lambda=0.107$, p-value = 0.05) — a meaningful shift toward less competitive behavior among the remaining airlines — providing evidence that mergers can have coordinated effects.
@@
Price Effects. Panel D presents the merger effects on prices, incorporating both efficiency and conduct channels. As with the previous estimates, none of the price effects is estimated precisely. DL-NW shows a price decrease of $\mathdollar 8.6$ per hundred dollars of fare, reflecting the combination of efficiency gain and reduced conduct. UA-CO exhibits a modest price increase of $\mathdollar 2.6$ per hundred dollars, while AA-US shows a larger price increase of $\mathdollar 10.6$ per hundred dollars. These effects reflect the combined impact of efficiency changes and altered competitive conduct following the merger.
The progressive deterioration in competitive conduct – from negative to positive to statistically significant positive – suggests that successive mergers may have facilitated increasingly coordinated behavior among the remaining carriers. Indeed, this escalating pattern of coordination effects was precisely the concern that led the US DOJ to initially block the AA-US merger in 2013, fearing it would lead to further coordinated effects. While we cannot make definitive causal claims due to statistical limitations, the systematic temporal progression provides suggestive evidence that merger waves in concentrated industries may have cumulative effects on market competitiveness.